OpenStudy (anonymous):

What is the simplified form of 5√x*5√x*5√x*5√x

OpenStudy (anonymous):

\[\sqrt{x}.\sqrt{x}=x\]

OpenStudy (anonymous):

5/x^2

OpenStudy (anonymous):

\[5\sqrt{x}~~~X~~5\sqrt{x}=?\]

OpenStudy (anonymous):

@inowalst